A group of researchers tested whether snakes tend to choose a warm resting site when both a warm site and a cool site are presented to them. Their hypotheses were: H0 = Snakes do not prefer the warmer site. And HA = Snakes prefer the warmer site. They carried out the experiment, and with their data calculated a one-tailed P-value of P = 0.03. If they were using an α value of 0.05, and instead used a two-tailed test, they would have:

failed to reject the null hypothesis

rejected the null hypothesis

accepted the alternative hypothesis

This is a situation where it is too close to call.
Answer #1

For a one tailed test we are given the p-value here is 0.03 < 0.05 which is the level of significance, therefore the test is significant here and we reject the null hypothesis here.

For a two tailed test, the p-value is computed here as:
= 2 * p-value for a one tailed test

= 2*0.03 = 0.06

As the p-value for the two tailed test is 0.06 > 0.05, therefore the test is not significant here and we cannot reject the null hypothesis here. Therefore failed to reject the null hypothesis is the correct answer here.